reese:

A hiker in Africa discovers a skull that contains 7% of its original amount of C-14 find the age of the skull to the nearest year

reese:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @reese A hiker in Africa discovers a skull that contains 7% of its original amount of C-14 find the age of the skull to the nearest year \(\color{#0cbb34}{\text{End of Quote}}\)

sillybilly123:

For exponential decay, with decay constant \(k\): \(C = C_o e^{- k t}\) For C-14 half life \(\tau (\approx 5730 \text{ years})\), we have : \(\dfrac{1}{2} = e^{- k \tau} \implies k = \dfrac{\ln 2}{\tau}\) For \(C = C_o \times 0.07\), we are solving: \(0.07 = \exp\left( - \dfrac{\ln 2}{5730} t\right) \implies t = 21, 983\).... "to the nearest year"..... Closed, so no rules broken ;)